# Multi-scale direction-aware SAR object detection network via global information fusion

### v4

@@ 4.3.4 Comparison of different FPN architecture
4.4 Comparisons with previous methods
4.4.1 Comparison results on SAR-AIRcraft-1.0
We use the SAR-AIRcraft-1.0 dataset to train and test several common benchmark detection methods, including two-stage detectors: Faster R-CNN [29], Cascade R-CNN [30], and one-stage detectors: YOLOv6 [23], YOLOv8 [31]. In addition to this, we also selected three latest SAR aircraft detection methods: SKG-Net [32], SA-Net [25] and MLSDNet [33], and some formidable model, HRLE-SARDet[34], RT-DETR [37], CCDN-DETR [38], FBR-Net [35], AFSar [36]. Most of the methods are also applied to SSDD and HRSID. Table 7 shows the detection results under different methods.
Table 8: Comparison with the latest object detection methods and SAR ship detection methods on SSDD.
@@ 4.4 Comparisons with previous methods
In the Faster R-CNN method, the accuracy is 77.2%, which shows to a certain extent that the SAR-AIRcraft-1.0 dataset has a certain detection difficulty because aircraft objects of different categories have similar structures and sizes, and the same category also presents multi-scale characteristics. As illustrated in Table 7, SAR-Net achieves the highest performance in nearly every category, primarily due to its integration of global information, and on this basis, we only add a small amount of computational overhead, but the performance has been greatly improved. Considering the advanced nature of YOLOv8, we also experimented with its variants from small to large versions. We observed that the performance fluctuated within a range of 0.3%, accompanied by a notable increase in computational cost. In addition, the detection accuracy of each category has certain differences. For example, A320/321 performs best in different algorithms compared to other categories. This is mainly due to the special size of A320/321, with a fuselage length of more than 40 meters, making it easy to distinguish. For some objects in other categories, such as ARJ21 and Boeing aircraft, due to their multi-scale characteristics and varied directions, this presents a challenge for the network in learning invariant features within the same category. We enhance the detection capabilities of multi-scale and directional information-rich objects by fusing global information and performing compensation in deep and shallow layers respectively, as well as learning object direction-aware information.
4.4.2 Comparison results on SSDD
Table 9: Comparison with the latest object detection methods and SAR ship detection methods on HRSID.
@@ 4.4.2 Comparison results on SSDD
As indicated in the data presented in Table 8, our method exhibits strong competitiveness even when transferred to ship tasks. It can be observed that due to the relatively small size of the SSDD dataset, the current methods have already achieved a high level of average precision. However, due to the continued presence of multi-scale objects and interference from background noise, this has prevented further performance improvements in other methods. Note that our method excels in information fusion and transmission, making it less prone to influence by these problems and less prone to overfit on small datasets. Consequently, this further elevates the benchmark performance of the dataset. Additionally, in comparison to the state-of-the-art (SOTA) model YOLOv8’s larger version, we still achieve a 0.4% precision advantage on mAP50, while reducing parameters and GFLOPS to one-fourth and one-sixth, respectively.
@@
4.4.3 Comparison results on HRSID
Table 10: Comparison with the latest oriented SAR ship object detection methods on SSDD and HRSID.
@@ 4.4.3 Comparison results on HRSID
Continuing our evaluation, we chose state-of-the-art object detection methods, some of which were specially designed for SAR imagery. Compared with other approaches, our method maintains excellent performance with a relatively small computational cost, as shown in Table 9. Consistent with the results from SSDD, when compared to YOLOv8m, our method exhibits a 0.6% precision advantage on mAP50 and significantly smaller parameter size and GFLOPs, being one-fourth and one-sixth of YOLOv8m, respectively.
@@
4.4.4 Comparison results on SSDD and HRSID with oriented bounding box detection methods
To further explore the direction-aware attention module for detecting rotated objects, we compared it with state-of-the-art SAR image rotation detection methods and conducted experiments on two ship datasets with rotation annotations. The results, as illustrated in Table 10, demonstrate that our method achieved the highest accuracy on both datasets. Despite not being specifically designed for rotation detection, the direction-aware attention module is still able to learn object feature information from two directions, thus assigning greater attention weights to the detected objects. Additionally, it retains this directional feature to provide perceptual information beneficial for detection.
4.5 Result visualization
@@ 5 Conclusion
Declaration of Competing Interest
No conflict of interest exists in the submission of this manuscript, and all authors approve the manuscript for publication. We would like to declare that the work described was original research that has not been published previously, and is not under consideration for publication elsewhere.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 62121001, Grant 62322117, Grant 62371365, and Grant U22B2014; and in part by the Young Elite Scientist Sponsorship Program through the China Association for Science and Technology under Grant 2020QNRC001.

### v5

@@ IV-B Accuracy Metrics
TABLE VIII: Comparison with the Latest Object Detection Methods and SAR Ship Detection Methods on SSDD
@@
IV-C Ablation Study
TABLE IX: Comparison with the Latest Object Detection Methods and SAR Ship Detection Methods on HRSID
@@ IV-C Ablation Study
IV-C1 Impact of Unity Compensation Mechanism
To study the effectiveness of each module in the unity compensation mechanism and verify the correctness of uniting deep and shallow layers’ global information, we executed a lot of ablation experiments on four of the modules and measured the performance with mAP50 and F1-score, a common indicator in object detection. In addition, Parameter sizes and GFLOPs are also shown in Table II and Table III. We use YOLOv6n [53] as the baseline for ablation experiments to embed the modules of our unity compensation mechanism one by one. The feature alignment process is included in all experiments except the baseline. In Table II, we conducted MAM using concatenation operation, where S-MFM represents Shallow-MFM and D represents Deep-MFM. It can be seen that only when the MAM module is used, the accuracy is lower than the baseline. This is because it integrates semantic information and detailed texture information from different layers, but it lacks further fusion of object-specific information. In addition, other experiments have achieved performance improvements without a substantial increase in computational overhead, which proves the efficacy of our global information fusion idea. In particular, when the MEM module is not used, we use simple addition to compensate for the original features.
To explore a new alignment pattern, we conducted MAM using addition operation. It is worth noting that both concatenation and addition operations require aligning features from each layer to have the same dimensions. As can be seen from Table III, the overall trends of the other modules are similar to those in Table II. However, due to the addition operation combining rich texture and semantic information directly from different layers, it made the information redundant and resulted in difficulties in subsequent fusion compensation operations, leading to information loss and performance degradation. Additionally, features from different layers have varying channel numbers, requiring channel alignment during the addition operation, which increases parameter and computation complexity.
IV-C2 Impact of Direction-aware Attention Module
To prove the effectiveness of the proposed direction-aware attention module, we executed a series of detailed experiments, as shown in Table IV. We split the attention modules in two directions and embedded them in the YOLOv6n backbone separately, both received performance improvements. To explore the roles of the two modules within DAM, we conducted experiments by using only one module at a time on the baseline. Here, w/o CAE represents without Channel Attention Embedding, and w/o DG represents without Direction-aware Generation. Both experiments showed improvement compared to the baseline. However, due to the absence of crucial bidirectional feature vector extraction, DAM w/o DG showed minimal performance enhancement, merely reallocating attention among channels. Equally noteworthy, after generating the direction-aware attention map, omitting the embedding of the attention map among channels also resulted in performance degradation. Therefore, the collaboration of these two modules enables the model to concentrate more effectively on the directional areas of objects.
@@ IV-C2 Impact of Direction-aware Attention Module
IV-C3 Collaboration of the Framework
The effectiveness of the components that make up SAR-Net has been demonstrated, and to further explore the synergy between them, we combine them and find the desired results. For the unity compensation mechanism, the backbone that can extract direction-aware information gives full play to the advantages of global fusion and enriches the information of multi-scale features. For the DAM-embedded backbone, the unity compensation mechanism retains detailed information and enhances the efficiency of information fusion and transmission. As shown in Table V, after combining the two, their collaborative complementarity is fully utilized and performance is further improved by 0.4%.
IV-C4 Comparison of Different FPN Architecture
We choose the classic FPN structure and its latest variants AFPN [49] and RepBi-PAN [53] for comparison. We conduct the experiments based on YOLOv6n and the results are illustrated in Table VI. A significant accuracy disparity exists despite the parameters being similar. As we discuss in the section I, sequential local fusion will lead to information loss, especially for SAR images with large object scale variants. Top-down and bottom-up structures will lose a lot of detailed information and semantic information before detection. Global information fusion and compensation preserve the object’s multi-scale information to the maximum extent, thereby enabling the unity compensation mechanism to deliver optimal performance.
IV-D Comparisons with Previous Methods
IV-D1 Comparison Results on SAR-AIRcraft-1.0
We use the SAR-AIRcraft-1.0 dataset to train and test several common benchmark detection methods, including two-stage detectors: Faster R-CNN [65], Cascade R-CNN [66], and one-stage detectors: YOLOv6 [53], YOLOv8 [67]. In addition to this, we also selected three latest SAR aircraft detection methods: SKG-Net [68], SA-Net [60] and MLSDNet [69], and some formidable model, HRLE-SARDet[70], RT-DETR [73], CCDN-DETR [74], FBR-Net [72], AFSar [71]. Most of the methods are also applied to SSDD and HRSID. Table VII shows the detection results under different methods.
TABLE X: Comparison with the Latest Oriented SAR Ship Object Detection Methods on SSDD and HRSID
@@ IV-D Comparisons with Previous Methods
In the Faster R-CNN method, the accuracy is 77.2%, which shows to a certain extent that the SAR-AIRcraft-1.0 dataset has a certain detection difficulty because aircraft objects of different categories have similar structures and sizes, and the same category also presents multi-scale characteristics. As illustrated in Table VII, SAR-Net achieves the highest performance in nearly every category, primarily due to its integration of global information, and on this basis, we only add a small amount of computational overhead, but the performance has been greatly improved. Considering the advanced nature of YOLOv8, we also experimented with its variants from small to large versions. We observed that the performance fluctuated within a range of 0.3%, accompanied by a notable increase in computational cost. In addition, the detection accuracy of each category has certain differences. For example, A320/321 performs best in different algorithms compared to other categories. This is mainly due to the special size of A320/321, with a fuselage length of more than 40 meters, making it easy to distinguish. For some objects in other categories, such as ARJ21 and Boeing aircraft, due to their multi-scale characteristics and varied directions, this presents a challenge for the network in learning invariant features within the same category. We enhance the detection capabilities of multi-scale and directional information-rich objects by fusing global information and performing compensation in deep and shallow layers respectively, as well as learning object direction-aware information.
IV-D2 Comparison Results on SSDD
As indicated in the data presented in Table VIII, our method exhibits strong competitiveness even when transferred to ship tasks. It can be observed that due to the relatively small size of the SSDD dataset, the current methods have already achieved a high level of average precision. However, due to the continued presence of multi-scale objects and interference from background noise, this has prevented further performance improvements in other methods. Note that our method excels in information fusion and transmission, making it less prone to influence by these problems and less prone to overfit on small datasets. Consequently, this further elevates the benchmark performance of the dataset. Additionally, in comparison to the state-of-the-art (SOTA) model YOLOv8’s larger version, we still achieve a 0.4% precision advantage on mAP50, while reducing parameters and GFLOPS to one-fourth and one-sixth, respectively.
@@ IV-D2 Comparison Results on SSDD
IV-D3 Comparison Results on HRSID
Continuing our evaluation, we chose state-of-the-art object detection methods, some of which were specially designed for SAR imagery. Compared with other approaches, our method maintains excellent performance with a relatively small computational cost. Consistent with the results from SSDD, when compared to YOLOv8m, our method exhibits a 0.6% precision advantage on mAP50 and significantly smaller parameter size and GFLOPs, being one-fourth and one-sixth of YOLOv8m, respectively.
IV-D4 Comparison Results on SSDD and HRSID with Oriented Bounding Box Detection Methods
To further explore the direction-aware attention module for detecting rotated objects, we compared it with state-of-the-art SAR image rotation detection methods and conducted experiments on two ship datasets with rotation annotations. The results, as illustrated in Table X, demonstrate that our method achieved the highest accuracy on both datasets. Despite not being specifically designed for rotation detection, the direction-aware attention module is still able to learn object feature information from two directions, thus assigning greater attention weights to the detected objects. Additionally, it retains this directional feature to provide perceptual information beneficial for detection.
IV-E Result Visualization